OpenStudy (anonymous):

y/-4.5 lesser than or = -10

OpenStudy (anonymous):

\[\frac{y}{-4.5}\le-10\] \[y \ge 45\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem!